Team A scored 5 more points than team B. Together, both teams scored a total of 151 points. how many points did team A score during this game?
Answer by waynest(281)

team A = x
team B = x + 5
total = 151
x + (x + 5) = 151
2x + 5 = 151
2x + 5 - 5 = 151 - 5
2x = 146
2x/2 = 146/2
x = 73
check:
73 + (73 + 5) = 151
73 + 78 = 151
151 = 151